Please help me solve this age problem: Carol is now thrice as old as Camille. Carols age three years ago minus the age of Camille in 5 years is 8. How old are they now?
Answer by TimothyLamb(4379)

x = carol age now
y = camille age now
---
x = 3y
x - 3 - (y + 5) = 8
---
put the system of linear equations into standard form
---
x - 3y = 0
x - 3 - y - 5 = 8
---
x - 3y = 0
x - y = 16
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = carol age now = 24
y = camille age now = 8
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php